# Universally Optimal Periodic Configurations in the Plane

### v3

@@ 5.1 $\tilde{F}\geq\tilde{g}$ on $[-1,1]\times([-1,-1/2]\cup[1/2,1])$
For fixed $t_{2}$, $\tilde{F}(t_{1},t_{2})-\tilde{g}(t_{1},t_{2})$ is strictly convex on $[-1,1]$ as a function of $t_{1}$ since $\tilde{g}(t_{1},t_{2})$ is linear in $t_{1}$ and $\tilde{f_{1}}$ is strictly absolutely monotone. The next proposition is an immediate consequence of this observation.
@@
We have $\tilde{F}\geq\tilde{g}$ on $[-1,1]\times[1/2,1]$ with equality only at $(-1,1/2)$.
@@
We first note that $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(t_{1},t_{2})=\tilde{f_{1}}^{\prime}(t_{1})\tilde{f_{2}}(t_{2})-a_{1,0}-a_{0,2}t_{2}$ is (a) strictly increasing in $t_{1}$ for fixed $t_{2}$ and (b) strictly convex in $t_{2}$ for fixed $t_{1}$. Let $h(t_{2}):=\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(-1,t_{2})$. The inequality (72) together with (a) implies $h(-1/2)<0$. Hence, the strict convexity of $h$ together with $h(1/2)>0$ (from (71)) implies $h(t_{2})=\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(-1,t_{2})>0$ for $t_{2}\in[1/2,1].$ Combining this fact with Lemma 23, we may invoke Lemma 26 part (a) to complete the proof. ∎
@@
Suppose by way of contradiction that there exists $t_{2}^{\prime}\in[-1,1]$ such that $t_{2}^{\prime}\neq-\frac{1}{2}$ and $\tilde{F}(1,t_{2}^{\prime})\leq\tilde{g}(1,t_{2}^{\prime})$. Then there must be some point $-1/2\neq p\in[-1,1]$ such that $\tilde{f_{1}}(1)\tilde{f_{2}}(p)=\tilde{g}(1,p)$. Indeed, either $t_{2}^{\prime}$ is such a point, or $\tilde{F}(1,t_{2}^{\prime})<\tilde{g}(1,t_{2}^{\prime})$. We have from Lemmas 25, 26, and 27 that $\tilde{F}(1,\pm 1)>\tilde{g}(1,\pm 1)$, which yield two cases for $t_{2}^{\prime}$. If $t_{2}^{\prime}<-1/2$, then there exists $p\in(-1,t_{2}^{\prime})$ such that $\tilde{F}(1,p)=\tilde{g}(1,p)$ by the intermediate value theorem. If $t_{2}^{\prime}>-1/2$, instead apply the intermediate value theorem on the interval $[t_{2}^{\prime},1]$ to see that $p\in[1/2,1]$.
@@
We have $\tilde{F}\geq\tilde{g}$ on $[-1,1]\times[-1,-1/2]$ with equality only at $(-1,-1)$ and $(1,-1/2)$.
@@
From Lemma 23, we have $\tilde{F}\geq\tilde{g}$ for $t_{1}=-1$. By Lemmas 25 and 26, we have the same inequality when $t_{2}=-1/2$ or $-1$. Finally, by Lemma 28, we have the inequality for $t_{1}=1$. All of these inequalities are strict except for at $(-1,-1)$ and $(1,-1/2)$.
Let $p=(p_{1},p_{2})$ be an arbitrary point on the boundary of $[-1,1]\times[-1,-1/2]$ such that $p_{1}<1$ and $p_{2}<-1/2$, let $q=(1,-1/2)$, and let $l(s):=p+s(q-p)$, $0\leq s\leq 1$, parametrize the line segment from $p$ to $q$. Since $u_{l}:=q-p$ has positive components, it follows that $\tilde{F}^{l}:=\tilde{F}\circ l$ is strictly absolutely monotone on $[0,1]$. Also, let $\tilde{g}^{l}:=\tilde{g}\circ l$ and note that $\tilde{g}^{l}$ is a polynomial of degree at most 2.
We claim that $(\tilde{F}^{l}-\tilde{g}^{l})(\epsilon)>0$ for all sufficiently small $\epsilon>0$. Indeed, if $p\neq(-1,-1)$, then $(\tilde{F}^{l}-\tilde{g}^{l})(0)>0$ and the result follows by continuity. If $p=(-1,-1)$, then $\nabla(\tilde{F}-\tilde{g})(-1,-1)\cdot u_{l}>0$ at (-1,-1) by Lemmas 25 and 23 which shows the result in this case. Similarly, the necessary derivative inequality and equality conditions at $(1,-1/2)$ imply $\nabla(\tilde{F}-\tilde{g})(-1,1/2)\cdot u_{l}<0$. Together with the fact that $(\tilde{F}^{l}-\tilde{g}^{l})(1)=0$, we get $(\tilde{F}^{l}-\tilde{g}^{l})(1-\epsilon)>0$ for all $\epsilon$ sufficiently small.
Now supposing for a contradiction that $(\tilde{F}^{l}-\tilde{g}^{l})(r^{\prime})<0$ for some $r^{\prime}\in(0,1)$. By the intermediate value theorem there are points $10<r_{1}<r^{\prime}<r_{2}<$ 1 such that $(\tilde{F}^{l}-\tilde{g}^{l})(r_{1})=(\tilde{F}^{l}-\tilde{g}^{l})(r_{2})=0$. Then $\tilde{g}^{l}$ is a polynomial of degree at most 2 which interpolates $\tilde{F}^{l}$ for $T=\{r_{1},r_{2},1\}$ and leads to a contradiction using the error formula (81). Since any point in $(-1,1)\times(-1,-1/2)$ must lie on such a line segment, we conclude that $\tilde{F}\geq\tilde{g}$ on $(-1,1)\times(-1,-1/2)$. Now to see the inequality must be strict, if $(\tilde{F}^{l}-\tilde{g}^{l})(r^{\prime})=0$ for some $r^{\prime}\in(0,1)$ $\tilde{g}^{l}$ is a polynomial of degree at most 2 which interpolates $\tilde{F}^{l}$ for $T=\{r^{\prime},r^{\prime},1\}$, and again we obtain a contradiction with the error formula. ∎
Thus, we have proved $\tilde{F}\geq\tilde{g}$ on $[-1,1]^{2}$ whenever $t_{2}\geq 1/2$ or $t_{2}\leq-1/2$. Our proof of the inequality for the critical region $-1/2\leq t_{2}\leq 1/2$ is more delicate and requires different approaches for $a$ small and $a$ large.
@@ 5.2 The critical region for small $a$ ($a<\pi^{2}$)
where the second inequality uses that $L_{\pm 1}(t_{1},t_{2})$ is a linear polynomial in $t_{1}$ for fixed $t_{2}$. Note the first inequality in (74) is strict if $-1<t_{1}<1$.
@@
where $-1<\chi_{-}<0<\chi_{+}<1$ which with (75) and the absolute monotonicity of $\tilde{f_{1}}$ implies
@@
then (75) and Lemma 74 show $(\tilde{F}-\tilde{g})(t_{1},t_{2})>0$ for all $-1<t_{1}<1.$ So it suffices to show $\phi\geq 0$ on $[-1/2,1/2]$ to prove that $\tilde{g}\leq\tilde{F}$ on the critical region. We can express $\phi(t_{2})$ as
@@
It follows from Lemma 32 that $A+B(t_{2}+1/2)+\frac{C}{2}(t_{2}+1/2)^{2}>0$ for $t_{2}\geq-1/2$ completing the proof that $\tilde{g}\leq\tilde{F}$ in the case $a<\pi^{2}$, and moreover, showing that $\tilde{F}=\tilde{g}$ only at our interpolation points $(-1,1),(-1,1/2),(1,-1/2)$.
@@
We have $\tilde{F}\geq\tilde{g}$ on $A=[-1,-\sqrt{2}/2]\times[0,1/2]$ with equality only at $(-1,1/2)$.
@@
First, we show the inequality for $[-1,-\sqrt{2}/2]\times[1/4,1/2]$. Since we already have $\tilde{F}\geq\tilde{g}$ when $t_{2}=1/2$ or $t_{1}=-1$, it suffices by Lemma 20 to show inequalities (62) on the 2 segments when $t_{2}=1/4$ or $t_{1}=-\sqrt{2}/2$, which we handle in the appendix Section D.4.3. Now having $\tilde{F}\geq\tilde{g}$ on the segment of $[-1,-\sqrt{2}/2]\times[0,1/4]$ when $t_{2}=1/4$, we again show inequalities (62) with $\tilde{g}_{-1,1/4}$ on the segments when $t_{2}=0$ or $t_{1}=-\sqrt{2}/2$ to complete the proof.
@@
We have $\tilde{F}\geq\tilde{g}$ on $C=[0,1]\times[-1/2,0]$ with equality only at $(1,-1/2)$.
@@
Thus, $\tilde{F}-\tilde{g}$ is minimized along the right and bottom boundaries of the region, where we have already showed $\tilde{F}-\tilde{g}\geq 0$ in the previous section with equality only at $(1,-1/2)$. The inequality $L_{1}(t_{1},t_{2})>0$ for $(t_{1},t_{2})\in[0,1]\times[-1/2,0]$ is proved in the appendix section D.4.5. ∎
@@
Lemma 37.
[…]
Proof.
We extend the domain of the $\log$ function so that $\log(t)=\infty$ for $t\leq 0$. Note that this definition and the fact that $\tilde{F}>0$ on all of $[-1,1]^{2}$ imply $\log(\tilde{F}/\tilde{g})>0$ on $E$ is equivalent to $\tilde{F}>\tilde{g}$ on $E$. Since we have already established that this inequality holds on $\partial E$, it suffices to show $\log(\tilde{F}/\tilde{g})$ takes no finite local minima on $S$, which we’ll do by showing that
@@
[…]
@@
Acknowledgements. The authors thank Henry Cohn, Denali Relles, Larry Rolen, Ed Saff, and Yujian Su for helpful discussions at various stages of this project.
@@ Appendix A Divided differences and univariate interpolation
We review some basic results concerning one-dimensional polynomial interpolation (e.g., see [4, Section 5.6.2]). Let $f\in C^{m}[a,b]$ for some $a,b$ be given along with some multiset
@@
Then there then exists a unique polynomial $H_{T}(f)(t)$ of degree at most $m$ (called a Hermite interpolant of $f$) such that for each $\alpha\in T$, we have $H^{(\ell)}_{T}(f)(\alpha)=f^{(\ell)}(\alpha)$ for $0\leq\ell<k_{\alpha}$ where $k_{\alpha}$ denotes the multiplicity of $\alpha$ in $T$. Let $f[t_{0},...,t_{m}]$ denote the coefficient of $t^{m}$ in $H_{T}(f)(t)$. This coefficient is called the m-th divided difference of $f$ for $T$. Then $H_{T}(f)$, may be expressed as
@@ Appendix B Equivalence of Different Notions of Universal Optimality
which is proved analogously to [4, Theorem 8.4.1].

### v4

@@ 5.1 $\tilde{F}\geq\tilde{g}$ on $1,1]\times([-1,-1/2]\cup[1/2,1])$
For fixed $t_{2}$, $\tilde{F}(t_{1},t_{2})-\tilde{g}(t_{1},t_{2})$ is strictly convex on [1,1][-1,1] as a function of $t_{1}$ since $\tilde{g}(t_{1},t_{2})$ is linear in $t_{1}$ and $\tilde{f_{1}}$ is strictly absolutely monotone. The next proposition is an immediate consequence of this observation.
@@
We have $\tilde{F}\geq\tilde{g}$ on $1,1]\times[1/2,1]$ with equality only at (1,1/2)(-1,1/2).
@@
We first note that $\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(t_{1},t_{2})=\tilde{f_{1}}^{\prime}(t_{1})\tilde{f_{2}}(t_{2})-a_{1,0}-a_{0,2}t_{2}$ is (a) strictly increasing in $t_{1}$ for fixed $t_{2}$ and (b) strictly convex in $t_{2}$ for fixed $t_{1}$. Let $h(t_{2}):=\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(-1,t_{2})$. The inequality (72) together with (a) implies h(1/2)<0h(-1/2)<0. Hence, the strict convexity of $h$ together with $h(1/2)>0$ (from (71)) implies $h(t_{2})=\frac{\partial(\tilde{F}-\tilde{g})}{\partial t_{1}}(-1,t_{2})>0$ for $t_{2}\in[1/2,1].$ Combining this fact with Lemma 23, we may invoke Lemma 26 part (a) to complete the proof. ∎
@@
Suppose by way of contradiction that there exists $t_{2}^{\prime}\in[-1,1]$ such that $t_{2}^{\prime}\neq-\frac{1}{2}$ and $\tilde{F}(1,t_{2}^{\prime})\leq\tilde{g}(1,t_{2}^{\prime})$. Then there must be some point $1/2\neq p\in[-1,1]$ such that $\tilde{f_{1}}(1)\tilde{f_{2}}(p)=\tilde{g}(1,p)$. Indeed, either $t_{2}^{\prime}$ is such a point, or $\tilde{F}(1,t_{2}^{\prime})<\tilde{g}(1,t_{2}^{\prime})$. We have from Lemmas 25, 26, and 27 that $\tilde{F}(1,\pm 1)>\tilde{g}(1,\pm 1)$, which yield two cases for $t_{2}^{\prime}$. If $t_{2}^{\prime}<-1/2$, then there exists $p\in(-1,t_{2}^{\prime})$ such that $\tilde{F}(1,p)=\tilde{g}(1,p)$ by the intermediate value theorem. If $t_{2}^{\prime}>-1/2$, instead apply the intermediate value theorem on the interval $t_{2}^{\prime},1]$ to see that $p\in[1/2,1]$.
@@
We have $\tilde{F}\geq\tilde{g}$ on $1,1]\times[-1,-1/2]$ with equality only at (1,1)(-1,-1) and (1,1/2)(1,-1/2).
@@
From Lemma 23, we have $\tilde{F}\geq\tilde{g}$ for $t_{1}=-1$. By Lemmas 25 and 26, we have the same inequality when $t_{2}=-1/2$ or 1-1. Finally, by Lemma 28, we have the inequality for $t_{1}=1$. All of these inequalities are strict except for at (1,1)(-1,-1) and (1,1/2)(1,-1/2).
Let $p=(p_{1},p_{2})$ be an arbitrary point on the boundary of $1,1]\times[-1,-1/2]$ such that $p_{1}<1$ and $p_{2}<-1/2$, let q=(1,1/2)q=(1,-1/2), and let l(s):=p+s(qp)l(s):=p+s(q-p), $0\leq s\leq 1$, parametrize the line segment from $p$ to $q$. Since $u_{l}:=q-p$ has positive components, it follows that $\tilde{F}^{l}:=\tilde{F}\circ l$ is strictly absolutely monotone on $[0,1]$. Also, let $\tilde{g}^{l}:=\tilde{g}\circ l$ and note that $\tilde{g}^{l}$ is a polynomial of degree at most 2.
We claim that $(\tilde{F}^{l}-\tilde{g}^{l})(\epsilon)>0$ for all sufficiently small $\epsilon>0$. Indeed, if $p\neq(-1,-1)$, then $(\tilde{F}^{l}-\tilde{g}^{l})(0)>0$ and the result follows by continuity. If p=(1,1)p=(-1,-1), then $\nabla(\tilde{F}-\tilde{g})(-1,-1)\cdot u_{l}>0$ at (-1,-1) by Lemmas 25 and 23 which shows the result in this case. Similarly, the necessary derivative inequality and equality conditions at (1,1/2)(1,-1/2) imply $\nabla(\tilde{F}-\tilde{g})(-1,1/2)\cdot u_{l}<0$. Together with the fact that $(\tilde{F}^{l}-\tilde{g}^{l})(1)=0$, we get $(\tilde{F}^{l}-\tilde{g}^{l})(1-\epsilon)>0$ for all $\epsilon$ sufficiently small.
Now supposing for a contradiction that $(\tilde{F}^{l}-\tilde{g}^{l})(r^{\prime})<0$ for some $r^{\prime}\in(0,1)$. By the intermediate value theorem there are points $0<r_{1}<r^{\prime}<r_{2}<1$ such that $(\tilde{F}^{l}-\tilde{g}^{l})(r_{1})=(\tilde{F}^{l}-\tilde{g}^{l})(r_{2})=0$. Then $\tilde{g}^{l}$ is a polynomial of degree at most 2 which interpolates $\tilde{F}^{l}$ for $T=\{r_{1},r_{2},1\}$ and leads to a contradiction using the error formula (81). Since any point in $(-1,1)\times(-1,-1/2)$ must lie on such a line segment, we conclude that $\tilde{F}\geq\tilde{g}$ on $(-1,1)\times(-1,-1/2)$. Now to see the inequality must be strict, if $(\tilde{F}^{l}-\tilde{g}^{l})(r^{\prime})=0$ for some $r^{\prime}\in(0,1)$ $\tilde{g}^{l}$ is a polynomial of degree at most 2 which interpolates $\tilde{F}^{l}$ for $T=\{r^{\prime},r^{\prime},1\}$, and again we obtain a contradiction with the error formula. ∎
Thus, we have proved $\tilde{F}\geq\tilde{g}$ on $1,1]^{2}$ whenever $t_{2}\geq 1/2$ or $t_{2}\leq-1/2$. Our proof of the inequality for the critical region $1/2\leq t_{2}\leq 1/2$ is more delicate and requires different approaches for $a$ small and $a$ large.
@@ 5.2 The critical region for small $a$ ($a<\pi^{2}$)
where the second inequality uses that $L_{\pm 1}(t_{1},t_{2})$ is a linear polynomial in $t_{1}$ for fixed $t_{2}$. Note the first inequality in (74) is strict if $1<t_{1}<1$.
@@
where $1<\chi_{-}<0<\chi_{+}<1$ which with (75) and the absolute monotonicity of $\tilde{f_{1}}$ implies
@@
then (75) and Lemma 74 show $(\tilde{F}-\tilde{g})(t_{1},t_{2})>0$ for all $1<t_{1}<1.$ So it suffices to show $\phi\geq 0$ on [1/2,1/2][-1/2,1/2] to prove that $\tilde{g}\leq\tilde{F}$ on the critical region. We can express $\phi(t_{2})$ as
@@
It follows from Lemma 32 that $A+B(t_{2}+1/2)+\frac{C}{2}(t_{2}+1/2)^{2}>0$ for $t_{2}\geq-1/2$ completing the proof that $\tilde{g}\leq\tilde{F}$ in the case $a<\pi^{2}$, and moreover, showing that $\tilde{F}=\tilde{g}$ only at our interpolation points (1,1),(1,1/2),(1,1/2)(-1,1),(-1,1/2),(1,-1/2).
@@
We have $\tilde{F}\geq\tilde{g}$ on $A=[-1,-\sqrt{2}/2]\times[0,1/2]$ with equality only at (1,1/2)(-1,1/2).
@@
First, we show the inequality for $1,-\sqrt{2}/2]\times[1/4,1/2]$. Since we already have $\tilde{F}\geq\tilde{g}$ when $t_{2}=1/2$ or $t_{1}=-1$, it suffices by Lemma 20 to show inequalities (62) on the 2 segments when $t_{2}=1/4$ or $t_{1}=-\sqrt{2}/2$, which we handle in the appendix Section D.4.3. Now having $\tilde{F}\geq\tilde{g}$ on the segment of $1,-\sqrt{2}/2]\times[0,1/4]$ when $t_{2}=1/4$, we again show inequalities (62) with $\tilde{g}_{-1,1/4}$ on the segments when $t_{2}=0$ or $t_{1}=-\sqrt{2}/2$ to complete the proof.
@@
We have $\tilde{F}\geq\tilde{g}$ on $C=[0,1]\times[-1/2,0]$ with equality only at (1,1/2)(1,-1/2).
@@
Thus, $\tilde{F}-\tilde{g}$ is minimized along the right and bottom boundaries of the region, where we have already showed $\tilde{F}-\tilde{g}\geq 0$ in the previous section with equality only at (1,1/2)(1,-1/2). The inequality $L_{1}(t_{1},t_{2})>0$ for $(t_{1},t_{2})\in[0,1]\times[-1/2,0]$ is proved in the appendix section D.4.5. ∎
@@
We first show inequalities (62) hold for $\tilde{g}_{-\sqrt{2}/2,0}$ on each line segment on the boundary of $\sqrt{2}/2,0\times[-.1,0]$ except the $t_{2}=0$ segment (where we already have $\tilde{F}>\tilde{g}$). Then we repeat the process with $\tilde{g}_{-\sqrt{2}/2,-.1}$ on each segment of $\sqrt{2}/2,0]\times[-.2,-.1]$ except the $t_{2}=-.1$ segment. The precise calculations are carried out in the appendix section D.4.3.
Lemma 37.
We have $\tilde{F}>\tilde{g}$ on $E=[-1,0]\times[-.5,-.2]\cup[-1,-\sqrt{2}/2]\times[-.5,0]$.
Proof.
@@
We will next establish that on all of $E$, $L_{2}$ is decreasing in $t_{1}$ and $t_{2}$. Thus to show $L_{2}>0$ on all of $E$, we need only check that $L_{2}(-\sqrt{2}/2,0),L_{2}(0,-.2)>0$, which is handled in the appendix section D.4.6. To see that $L_{2}$ is decreasing in both $t_{1}$ and $t_{2}$, observe by Proposition 10 that
@@
Appendix
@@ Appendix A Divided differences and univariate interpolation
We review some basic results concerning one-dimensional polynomial interpolation (e.g., see [3, Section 5.6.2]). Let $f\in C^{m}[a,b]$ for some $a,b$ be given along with some multiset
@@
Then there exists a unique polynomial $H_{T}(f)(t)$ of degree at most $m$ (called a Hermite interpolant of $f$) such that for each $\alpha\in T$, we have $H^{(\ell)}_{T}(f)(\alpha)=f^{(\ell)}(\alpha)$ for $0\leq\ell<k_{\alpha}$ where $k_{\alpha}$ denotes the multiplicity of $\alpha$ in $T$. Let $f[t_{0},...,t_{m}]$ denote the coefficient of $t^{m}$ in $H_{T}(f)(t)$. This coefficient is called the m-th divided difference of $f$ for $T$. Then $H_{T}(f)$, may be expressed as
@@ Appendix B Equivalence of Different Notions of Universal Optimality
which is proved analogously to [3, Theorem 8.4.1].